Find an equation of the line. Write the equation using function notation.

Through (−5,−9)​; parallel to 6x+5y=11  


1
Expert's answer
2022-02-15T17:17:25-0500

Solution

Re-write 6x+5y =11 into the form y=mx + c

5y=-6x+11

y = -6/5 x +11/5

Parallel lines have equal gradient m = -6/5


Hence, the equation of the line would be given by

"{y-(-9) \above{2pt} x-(-5)}=- {6 \above{2pt} 5}\\

5(y + 9) =-6(x + 5)\\
5y +45 = -6x-30\\
5y +6x = -75 [Answer]"